OpenStudy (anonymous):

How to solve 2p=7+q, 6p-3q=24

OpenStudy (anonymous):

combine like terms

OpenStudy (anonymous):

2p=7+q ->p=7/2+q/2 substitute it on the 2nd equation 6(7/2+q/2)-3q=24 =>21+3q-3q owwwwwwwww it has got no solution